What is the delta of a short position in 600 European call options on silver futures? The options mature in 8 months and the silver futures contract matures in 9 months. The current 9 month futures price is $30.00 per ounce. The exercise price of the option is $31.00 per ounce. The risk-free interest rate is 5% per year and the volatility is 20 percent per year.
Answer #1

Delta is defined as rate of change of derivative's price compared to that of underlying.

Give facts

Option Period= 8 months

Silver future contract period= 9 months

Futures Price( S)= 30

Exercise Price (X)= 31

Risk Free Rate (Rf)= 5%

Volatility= 20%

Delta= e-rt*N(d1)

d1= (in(s/x) + (r + σ2/2)t)/OVE

= (In(30/31) (.05 .20/2).6667)/.20V0.6667

=0.013869/0.1633

=0.0849

N(d1)= 0.5338

and Delta of Option is

.0 е *0.6667(0.5338)

0.9672*0.5338

=0.5163

The delta of a short position in 600 futures options is therefore= 309.78